OpenStudy (anonymous):

A plane surface having an area S and inclined at an angle B to the vertical is subjected to water pressure on one side only, the centroid being at a depth of H below the water surface. Working on first principles, obtain an expression giving the depth of the centre of pressure.

OpenStudy (mikurout):

= H+( square of (sinB) ) * \[\frac{ I _{G XX} }{ H \times S }\]